# Enhancing Variable Selection in Large-scale Logistic Regression: Leveraging Manual Labeling with Beneficial Noise

### v1

@@ IV-A ADMM for ALASSO-PLR with Manual Labels
ADMM is an algorithm widely used in convex optimization problems and can be applied to solve large-scale convex optimization problems with separable structures. This algorithm decomposes complex optimization problems into multiple simple subproblems and solves these subproblems through alternating iterations to gradually approach the optimal solution of the original problem. The augmented Lagrangian form of (IV) is
$L_{\mu}(\bm{\beta},\bm{r},\bm{u})$ $=\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\right]$
$+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\bm{u}^{\top}(\bm{X}\bm{\beta}-\bm{r})$ (21)
$+\frac{\mu}{2}\|\bm{X}\bm{\beta}-\bm{r}\|_{2}^{2},$
@@
where $\bm{u}$ is dual variable corresponding to the linear constraint, and $\mu>0$ is a given augmented parameter. The iterative scheme of ADMM for (IV) is
@@
$\left\{\begin{array}[]{l}\bm{\beta}^{k+1}\leftarrow\mathop{\arg\min}\limits_{\bm{\beta}}\left\{L_{\mu}(\bm{\beta},\bm{r}^{k},\bm{u}^{k})\right\};\\ \bm{r}^{k+1}\leftarrow\mathop{\arg\min}\limits_{\bm{r}}\left\{L_{\mu}(\bm{\beta}^{k+1},\bm{r},\bm{u}^{k})\right\};\\ \bm{u}^{k+1}\leftarrow\bm{u}^{k}-\mu(\bm{X}\bm{\beta}^{k+1}-\bm{r}^{k+1}).\end{array}\right.$ (22)
@@
For the subproblem of updating $\bm{\beta}^{k+1}$ in (22), we can rearrange the optimization equation and omit constant terms irrelevant to the target variable $\bm{\beta}$, yielding the following iterative formula,
$\bm{\beta}^{k+1}\leftarrow\arg\min_{\bm{\beta}}$$\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}\right.$ (23)
$+$ $\left.\frac{\mu}{2}\|\bm{X}\bm{\beta}-\bm{r}^{k}-\bm{u}^{k}/\mu\|_{2}^{2}\right\}.$
@@
Clearly, the non-identity matrix before the quadratic term of $\bm{\beta}$ and $\|\bm{w}\odot\bm{\beta}\|_{1}$ prevent iterative formula in (23) from having a closed-form solution. While numerical methods like coordinate descent can solve it, they greatly increase the computational burden. We use the linearization method from [32, 33, 34, 35] to approximate this optimization problem and obtain a closed-form solution for $\bm{\beta}^{k+1}$. Specifically, we suggest adding a proximal term to the objective function in (23) and updating $\bm{\beta}^{k+1}$ with the following iterative formula,
$\bm{\beta}^{k+1}\leftarrow\arg\min_{\bm{\beta}}$$\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}+\frac{\mu}{2}\|\bm{X}\bm{\beta}-\bm{r}^{k}-\bm{u}^{k}/\mu\|_{2}^{2}\right.$
$+$ $\left.\frac{1}{2}\|\bm{\beta}-\bm{\beta}^{k}\|^{2}_{{\bm{E}}_{1}}\right\}.$ (24)
@@
where $\bm{E}_{1}=\eta\bm{I}_{p}-\mu\bm{X}^{\top}\bm{X}$ is a positive semidefinite matrix. To guarantee that $\bm{E}_{1}$ is a positive definite matrix, $\eta$ should be no less than the largest eigenvalue of $\mu\bm{X}^{\top}\bm{X}$ (denoted as $\|\mu\bm{X}^{\top}\bm{X}\|$). After rearranging the terms in (IV-A), the update rule for $\bm{\beta}^{k+1}$ can be expressed as
$\bm{\beta}^{k+1}\leftarrow$ $\arg\min_{\bm{\beta}}\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}\right.$ (25)
$+$ $\left.\frac{\eta}{2}\|\bm{\beta}-\bm{\beta}^{k}+\frac{\mu\bm{X}^{\top}(\bm{X}\bm{\beta}^{k}-\bm{r}^{k}-\bm{u}^{k}/\mu)}{\eta}\|_{2}^{2}\right\}.$
@@
As a result, the closed-form solution of (25) is a weighted soft-thresholding operator, that is,
@@
$\bm{\beta}^{k+1}\leftarrow\text{sign}\left(\bm{\beta}^{k}-\frac{\mu\bm{X}^{\top}(\bm{X}\bm{\beta}^{k}-\bm{r}^{k}-\bm{u}^{k}/\mu)}{\eta}\right)$ (26)
$\odot\left(\left|\bm{\beta}^{k}-\frac{\mu\bm{X}^{\top}(\bm{X}\bm{\beta}^{k}-\bm{r}^{k}-\bm{u}^{k}/\mu)}{\eta}\right|-\lambda\bm{w}/\eta\right)_{+}.$
For the subproblem of updating $\bm{r}^{k+1}$ in (22), rearranging the optimization formula yields
$\bm{r}^{k+1}\leftarrow\arg\min_{\bm{r}}$$\left\{\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\right]\right.$
$+$ $\left.\frac{\mu}{2}\|\bm{X}\bm{\beta}^{k+1}-\bm{r}-\bm{u}^{k}/\mu\|_{2}^{2}\right\}.$ (27)
@@
Due to the existence of $m\log\left(1+\exp({r_{i}})\right)$, there is no closed-form solution for the above iterative formula. Linearization techniques can still be used, which involves adding a quadratic term, that is,
$\bm{r}^{k+1}$ $\leftarrow\arg\min_{\bm{r}}\left\{\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\right]\right.$ (28)
$+$ $\left.\frac{\mu}{2}\|\bm{X}\bm{\beta}^{k+1}-\bm{r}-\bm{u}^{k}/\mu\|_{2}^{2}+\frac{1}{2}\|\bm{r}-\bm{r}^{k}\|^{2}_{{\bm{E}}_{2}}\right\},$
@@
where $\bm{E}_{2}=(c_{0}-o(1))\bm{I}_{p}$ with $c_{0}$ is an arbitrary positive constant. Here, the $o(1)$ is only used for describing the algorithm iteration process, and there is no need to know its specific value. Adding this quadratic term is to remove the remainder of the second-order Taylor expansion of $\phi(r_{i})=m\log\left(1+\exp({r_{i}})\right)$. To be specific, for each $i$, we have
$\phi(r_{i})+(1-o(1))(r_{i}-r_{i}^{k})^{2}$
=\ = $\phi(r_{i}^{k})+\phi^{{}^{\prime}}(r_{i}^{k})(r_{i}-r_{i}^{k})$
+\ + $\frac{\phi^{{}^{\prime\prime}}(r_{i}^{k})+o(1)}{2}(r_{i}-r_{i}^{k})^{2}+\frac{c_{0}-o(1)}{2}(r_{i}-r_{i}^{k})^{2}$
=\ = $\phi(r_{i}^{k})+\phi^{{}^{\prime}}(r_{i}^{k})(r_{i}-r_{i}^{k})+\frac{\phi^{{}^{\prime\prime}}(r_{i}^{k})+c_{0}}{2}(r_{i}-r_{i}^{k})^{2}.$
Then, by rearranging the optimization equation (28), we obtain
$\bm{r}^{k+1}\leftarrow\arg\min_{\bm{r}}$$\left\{(\phi^{{}^{\prime}}(\bm{r}^{k})-\bm{S})^{\top}\bm{r}+\frac{1}{2}\|\bm{r}-\bm{r}^{k}\|^{2}_{\Phi(\bm{r}^{k})}\right.$
$+$ $\left.\frac{\mu}{2}\|\bm{X}\bm{\beta}^{k+1}-\bm{r}-\bm{u}^{k}/\mu\|_{2}^{2}\right\},$ (29)
@@
where $\phi^{{}^{\prime}}(\bm{r}^{k})=(\phi^{{}^{\prime}}(r_{1}^{k}),\phi^{{}^{\prime}}(r_{2}^{k}),\dots,\phi^{{}^{\prime}}(r_{n}^{k}))^{\top}$ and $\Phi(\bm{r}^{k})$ is a diagonal matrix, and the diagonal elements are $\{\phi^{{}^{\prime\prime}}(r_{i}^{k})+c_{0}\}_{i=1}^{n}$. Obviously, (IV-A) is a quadratic function, and the derivative can have the following closed-form solution,
$\bm{r}^{k+1}\leftarrow$ $\left[\Phi(\bm{r}^{k})+\mu\bm{I}_{p}\right]^{-1}\left[\mu(\bm{X}\bm{\beta}^{k+1}-\bm{u}^{k}/\mu)\right.$ (30)
$+$ $\left.\Phi(\bm{r}^{k})\bm{r}^{k}-\phi^{{}^{\prime}}(\bm{r}^{k})+\bm{S}\right].$
@@
To sum up, the iteration of ADMM for (IV) can be described in Algorithm 1. Note that Algorithm 1 can also be used to solve ALASSO-PLR under the truth label when $\bm{S}$ is generated by the real observation label $\bm{y}$.
@@
   1. Update $\bm{\beta}^{k+1}$ using (LABEL:lbeta3),
   2. Update $\bm{r}^{k+1}$ using (30),
@@
The design details of Algorithm 1 are entirely novel, even though the linearized ADMM algorithm has been a topic of long-standing discussion (for details, refer to Chapter 3 in [36] and the references therein). Specifically, the existing ADMM algorithms [31, 32, 33, 37] for PLR constructed equality constraints that replace $\bm{\beta}$ or its overlapping structure, while Algorithm 1 focuses on $\bm{X}\bm{\beta}=\bm{r}$. The use of constructing this equality constraint is convenient for the design of subsequent partitioning insensitive parallel algorithms. Moreover, Algorithm 1 linearizes both $\bm{\beta}$ and $\bm{r}$. This two-variable linearization approach has been reported in [33] (Algorithm 1) and [36] (Algorithms 3.2, 3.3). Both Algorithm 1 in [33] and Algorithm 3.2 in [36] focus on linearizing the quadratic augmented term. This stands in contrast to Algorithm 1 put forward in this paper. Specifically, in Algorithm 1 of this paper, the first term solely represents the linearized quadratic augmented term, while the second term is only related to the linearization of the logistic regression loss function. Algorithm 3.3 in [36] simultaneously linearizes both the loss function and the quadratic augmented term for two subproblems, which also distinguishes it from Algorithm 1 in this paper. Moreover, the linearization of its functions requires calculating the Lipschitz constant of the first-order derivative of some functions, differing from the linearization scheme proposed in this paper.
Next, we will discuss the convergence and convergence rate of Algorithm 1. Its proof can be found in Appendix A-C.
@@
    (Algorithm global convergence). It converges to some $\bm{v}^{\infty}=(\bm{\beta}^{\infty},\bm{r}^{\infty},\bm{u}^{\infty})$ that belongs to $\Omega^{*}$, where $\Omega^{*}$ is the set of optimal solutions for (IV).
@@
    $\|\bm{v}^{K}-\bm{v}^{K+1}\|_{\bm{H}}^{2}\leq\frac{1}{c_{0}\left(K+1\right)}\|\bm{v}^{0}-\bm{v}^{*}\|_{\bm{H}}^{2},$ (31)
    where $\bm{v}^{*}\in\Omega^{*}$, $c_{0}$ is a positive constant, and $\bm{H}=\begin{pmatrix}\bm{E}_{1}&0&0\\ 0&\mu\bm{I}_{n}+\bm{E}_{2}&0\\ 0&0&\frac{1}{\mu}\bm{I}_{n}\end{pmatrix}$ is a positive definite matrix.
@@
Although the differences between Algorithm 3.3 in [36] and Algorithm 1 have been discussed above, Algorithm 1 can be regarded as a special case of Algorithm 3.3 in terms of the linearization strategy. However, regarding the algorithm convergence, the convergence conclusion of Theorem 3 differs from that presented in Chapter 3.2.1 of [36]. To be specific, Theorem 3.7 in [36] discusses the sublinear convergence of the linearized algorithm. However, this sublinear convergence is concerned with the optimization objective function and constraints, while Theorem 3 in this paper focuses on the iterative point sequence. Moreover, in the course of proving this theorem, we showed that both $\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}$ and $\|\bm{v}^{k}-\bm{v}^{k+1}\|_{\bm{H}}^{2}$ exhibit non-increasing monotonicity. Specifically, we have $\|\bm{v}^{k+1}-\bm{v}^{*}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}$ and $\|\bm{v}^{k+1}-\bm{v}^{k+2}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{k+1}\|_{\bm{H}}^{2}$. For more details, refer to the propositions in the appendix.
@@
IV-B Parallel ADMM for ALASSO-PLR with Manual Labels

### v2

@@ 4.1 ADMM for ALASSO-PLR with Manual Labels
ADMM is an algorithm widely used in convex optimization problems (see [27, 28]) and can be applied to solve large-scale convex optimization problems with separable structures. This algorithm decomposes complex optimization problems into multiple simple subproblems and solves these subproblems through alternating iterations to gradually approach the optimal solution of the original problem. The augmented Lagrangian form of (4) is
$L_{\mu}(\bm{\beta},\bm{r},\bm{u})=\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\right]+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\bm{u}^{\top}(\bm{X}\bm{\beta}-\bm{r})+\frac{\mu}{2}\|\bm{X}\bm{\beta}-\bm{r}\|_{2}^{2},$ (4.2)
@@
where $\bm{u}$ is dual variable corresponding to the linear constraint, and $\mu>0$ is a given augmented parameter. The iterative scheme of ADMM for (4) is
@@
$\left\{\begin{array}[]{l}\bm{\beta}^{k+1}\leftarrow\mathop{\arg\min}\limits_{\bm{\beta}}\left\{L_{\mu}(\bm{\beta},\bm{r}^{k},\bm{u}^{k})\right\};\\ \bm{r}^{k+1}\leftarrow\mathop{\arg\min}\limits_{\bm{r}}\left\{L_{\mu}(\bm{\beta}^{k+1},\bm{r},\bm{u}^{k})\right\};\\ \bm{u}^{k+1}\leftarrow\bm{u}^{k}-\mu(\bm{X}\bm{\beta}^{k+1}-\bm{r}^{k+1}).\end{array}\right$. (4.3)
@@
For the subproblem of updating $\bm{\beta}^{k+1}$ in (4.3), we can rearrange the optimization equation and omit constant terms irrelevant to the target variable $\bm{\beta}$, yielding the following iterative formula,
$\bm{\beta}^{k+1}\leftarrow\arg\min_{\bm{\beta}}$ $\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}+\frac{\mu}{2}\|\bm{X}\bm{\beta}-\bm{r}^{k}-\bm{u}^{k}/\mu\|_{2}^{2}\right\}.$ (4.4)
@@
Clearly, the non-identity matrix before the quadratic term of $\bm{\beta}$ and $\|\bm{w}\odot\bm{\beta}\|_{1}$ prevent iterative formula in (4.4) from having a closed-form solution. While numerical methods like coordinate descent can solve it, they greatly increase the computational burden. We use the linearization method from [29] and [30] to approximate this optimization problem and obtain a closed-form solution for $\bm{\beta}^{k+1}$. Specifically, we suggest adding a proximal term to the objective function in (4.4) and updating $\bm{\beta}^{k+1}$ with the following iterative formula,
$\bm{\beta}^{k+1}\leftarrow\arg\min_{\bm{\beta}}\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}+\frac{\mu}{2}\|\bm{X}\bm{\beta}-\bm{r}^{k}-\bm{u}^{k}/\mu\|_{2}^{2}+\frac{1}{2}\|\bm{\beta}-\bm{\beta}^{k}\|^{2}_{{\bm{E}}_{1}}\right\}.$ (4.5)
@@
where $\bm{E}_{1}=\eta\bm{I}_{p}-\mu\bm{X}^{\top}\bm{X}$ is a positive semidefinite matrix. To guarantee that $\bm{E}_{1}$ is a positive definite matrix, $\eta$ should be no less than the largest eigenvalue of $\mu\bm{X}^{\top}\bm{X}$ (denoted as $\|\mu\bm{X}^{\top}\bm{X}\|$). After rearranging the terms in (4.5), the update rule for $\bm{\beta}^{k+1}$ can be expressed as
$\bm{\beta}^{k+1}\leftarrow\arg\min_{\bm{\beta}}\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}+\frac{\eta}{2}\|\bm{\beta}-\bm{\beta}^{k}+\frac{\mu\bm{X}^{\top}(\bm{X}\bm{\beta}^{k}-\bm{r}^{k}-\bm{u}^{k}/\mu)}{\eta}\|_{2}^{2}\right\}.$ (4.6)
@@
As a result, the closed-form solution of (4.6) is a weighted soft-thresholding operator, that is,
@@
$\bm{\beta}^{k+1}\leftarrow\text{sign}\left(\bm{\beta}^{k}-\frac{\mu\bm{X}^{\top}(\bm{X}\bm{\beta}^{k}-\bm{r}^{k}-\bm{u}^{k}/\mu)}{\eta}\right)\odot\left(\left|\bm{\beta}^{k}-\frac{\mu\bm{X}^{\top}(\bm{X}\bm{\beta}^{k}-\bm{r}^{k}-\bm{u}^{k}/\mu)}{\eta}\right|-\lambda\bm{w}/\eta\right)_{+}.$ (4.7)
For the subproblem of updating $\bm{r}^{k+1}$ in (4.3), rearranging the optimization formula yields
$\bm{r}^{k+1}\leftarrow\arg\min_{\bm{r}}\left\{\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\right]+\frac{\mu}{2}\|\bm{X}\bm{\beta}^{k+1}-\bm{r}-\bm{u}^{k}/\mu\|_{2}^{2}\right\}.$
@@
It is a system of nonlinear equations, and each component in the vector $\bm{r}$ can be solved element-by-element. Due to the existence of $m\log\left(1+\exp({r_{i}})\right)$, there is no closed-form solution for the above iterative formula. The Newton’s iteration method is an effective algorithm for solving this system of nonlinear equations. Linearization techniques can also be used, which involves adding a quadratic term, that is,
$\bm{r}^{k+1}\leftarrow\arg\min_{\bm{r}}\left\{\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\right]+\frac{\mu}{2}\|\bm{X}\bm{\beta}^{k+1}-\bm{r}-\bm{u}^{k}/\mu\|_{2}^{2}+\frac{1}{2}\|\bm{r}-\bm{r}^{k}\|^{2}_{{\bm{E}}_{2}}\right\},$ (4.8)
@@
where $\bm{E}_{2}=(c_{0}-o(1))\bm{I}_{p}$ with $c_{0}$ is an arbitrary positive constant. Note, when we use Newton’s method to solve the system of linear equations, setting $c_{0}=o(1)$ suffices. Here, the $o(1)$ is only used for describing the algorithm iteration process, and there is no need to know its specific value. Adding this quadratic term is to remove the remainder of the second-order Taylor expansion of $\phi(r_{i})=m\log\left(1+\exp({r_{i}})\right)$. To be specific, for each $i$, we have
$\phi(r_{i})+(1-o(1))(r_{i}-r_{i}^{k})^{2}$ $=\phi(r_{i}^{k})+\phi^{{}^{\prime}}(r_{i}^{k})(r_{i}-r_{i}^{k})+\frac{\phi^{{}^{\prime\prime}}(r_{i}^{k})+o(1)}{2}(r_{i}-r_{i}^{k})^{2}+\frac{c_{0}-o(1)}{2}(r_{i}-r_{i}^{k})^{2}$
$=\phi(r_{i}^{k})+\phi^{{}^{\prime}}(r_{i}^{k})(r_{i}-r_{i}^{k})+\frac{\phi^{{}^{\prime\prime}}(r_{i}^{k})+c_{0}}{2}(r_{i}-r_{i}^{k})^{2}.$
Then, by rearranging the optimization equation (4.8), we obtain
$\bm{r}^{k+1}\leftarrow\arg\min_{\bm{r}}\left\{(\phi^{{}^{\prime}}(\bm{r}^{k})-\bm{S})^{\top}\bm{r}+\frac{1}{2}\|\bm{r}-\bm{r}^{k}\|^{2}_{\Phi(\bm{r}^{k})}+\frac{\mu}{2}\|\bm{X}\bm{\beta}^{k+1}-\bm{r}-\bm{u}^{k}/\mu\|_{2}^{2}\right\},$ (4.9)
@@
where $\phi^{{}^{\prime}}(\bm{r}^{k})=(\phi^{{}^{\prime}}(r_{1}^{k}),\phi^{{}^{\prime}}(r_{2}^{k}),\dots,\phi^{{}^{\prime}}(r_{n}^{k}))^{\top}$ and $\Phi(\bm{r}^{k})$ is a diagonal matrix, and the diagonal elements are $\{\phi^{{}^{\prime\prime}}(r_{i}^{k})+c_{0}\}_{i=1}^{n}$. Obviously, (4.9) is a quadratic function, and the derivative can have the following closed-form solution,
$\bm{r}^{k+1}\leftarrow\left[\Phi(\bm{r}^{k})+\mu\bm{I}_{p}\right]^{-1}\left[\mu(\bm{X}\bm{\beta}^{k+1}-\bm{u}^{k}/\mu)+\Phi(\bm{r}^{k})\bm{r}^{k}-\phi^{{}^{\prime}}(\bm{r}^{k})+\bm{S}\right].$ (4.10)
@@
To sum up, the iteration of ADMM for (4) can be described in Algorithm 1. Note that Algorithm 1 can also be used to solve ALASSO-PLR under the truth label when $\bm{S}$ is generated by the real observation label $\bm{y}$.
@@
     1. Update $\bm{\beta}^{k+1}$ using (4.7),
     2. Update $\bm{r}^{k+1}$ using (4.10),
@@
The design details of Algorithm 1 are entirely novel, even though the linearized ADMM algorithm has been a topic of long-standing discussion (for details, refer to Chapter 3 in [31] and the references therein). Specifically, the existing ADMM algorithms in [29] and [32] for PLR constructed equality constraints that replace $\bm{\beta}$ or its overlapping structure, while Algorithm 1 focuses on $\bm{X}\bm{\beta}=\bm{r}$. The use of constructing this equality constraint is convenient for the design of subsequent partitioning insensitive parallel algorithms. Moreover, Algorithm 1 linearizes both $\bm{\beta}$ and $\bm{r}$. This two-variable linearization approach has been reported in [29] (Algorithm 1) and [31] (Algorithms 3.2, 3.3). Both Algorithm 1 in [29] and Algorithm 3.2 in [31] focus on linearizing the quadratic augmented term. This stands in contrast to Algorithm 1 put forward in this paper. Specifically, in Algorithm 1 of this paper, the first term solely represents the linearized quadratic augmented term, while the second term is only related to the linearization of the logistic regression loss function. Algorithm 3.3 in [31] simultaneously linearizes both the loss function and the quadratic augmented term for two subproblems, which also distinguishes it from Algorithm 1 in this paper. Moreover, the linearization of its functions requires calculating the Lipschitz constant of the first-order derivative of some functions, differing from the linearization scheme proposed in this paper.
Next, we will discuss the convergence and convergence rate of Algorithm 1. Its proof can be found in supplementary material 1.3.
@@
    (Algorithm global convergence). It converges to some $\bm{v}^{\infty}=(\bm{\beta}^{\infty},\bm{r}^{\infty},\bm{u}^{\infty})$ that belongs to $\Omega^{*}$, where $\Omega^{*}$ is the set of optimal solutions for (4).
@@
    $\|\bm{v}^{K}-\bm{v}^{K+1}\|_{\bm{H}}^{2}\leq\frac{1}{c_{0}\left(K+1\right)}\|\bm{v}^{0}-\bm{v}^{*}\|_{\bm{H}}^{2},$ (4.11)
    where $\bm{v}^{*}\in\Omega^{*}$, $c_{0}$ is a positive constant, $\|\bm{v}\|_{\bm{H}}=\sqrt{\bm{v}^{\top}\bm{H}\bm{v}}$ and $\bm{H}=\begin{pmatrix}\bm{E}_{1}&0&0\\ 0&\mu\bm{I}_{n}+\bm{E}_{2}&0\\ 0&0&\frac{1}{\mu}\bm{I}_{n}\end{pmatrix}$ is a positive definite matrix.
@@
Although the differences between Algorithm 3.3 in [31] and Algorithm 1 have been discussed above, Algorithm 1 can be regarded as a special case of Algorithm 3.3 in terms of the linearization strategy. However, regarding the algorithm convergence, the convergence conclusion of Theorem 3 differs from that presented in Chapter 3.2.1 of [31]. To be specific, Theorem 3.7 in [31] discusses the sublinear convergence of the linearized algorithm. However, this sublinear convergence is concerned with the optimization objective function and constraints, while Theorem 3 in this paper focuses on the iterative point sequence. Moreover, in the course of proving this theorem, we showed that both $\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}$ and $\|\bm{v}^{k}-\bm{v}^{k+1}\|_{\bm{H}}^{2}$ exhibit non-increasing monotonicity. Specifically, we have $\|\bm{v}^{k+1}-\bm{v}^{*}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}$ and $\|\bm{v}^{k+1}-\bm{v}^{k+2}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{k+1}\|_{\bm{H}}^{2}$. For more details, refer to the propositions in the supplementary material.
@@
4.2 Parallel ADMM for ALASSO-PLR with Manual Labels